# Quantum Magnetic Skyrmions on Kondo-type Lattices

### v3

@@ Abstract
A quantum description is given of nanoskyrmions in 2D textures with localised spins and itinerant electrons, isolated or coupled to leads, in or out-of- equilibrium. The spin-electron exchange is treated at the mean-field level, while Tensor Networks and exact diagonalization or nonequilibrium Green’s functions are used for localised spins and itinerant electrons. We motivate our scheme via exact and mean-field benchmarks, then show by several examples that itinerant electrons distinctly affect the properties of quantum nanoskyrmions. Finally, we mention lines of future work and improvement of the approach.
@@
Introduction. Magnetic skyrmions are topologically robust, long-lived spin textures which can be manipulated by ultra-low currents [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11], and of great potential for applications in spintronics and quantum computing [12, 13, 14, 15, 16]. Skyrmions originate from the interplay of different types of magnetic interactions in solids [17], such as Heisenberg exchange [18], Dzyaloshinskii-Moriya interaction (DMI) [19, 20], and anisotropy [21, 22], to mention some.
@@
Theoretical descriptions of skyrmions are often based on classical spin-only models [23, 24, 25, 26], treated via the Landau-Lifshitz-Gilbert [27, 28] and Thiele equations [29]. These formulations, where itinerant electrons (i-electrons) enter implicitly or as a source of renormalization of the classical spin dynamics, are in many cases adequate; however, describing explicitly the interaction (via Kondo-like exchange) between i-electrons and localised spins (l-spins) can also be important [30, 31, 32, 33, 34, 35, 36, 37, 38, 39], to e.g. address interface phenomena like current-driving and optical generation of nanoskyrmions [40].
@@
In most of currently available i-electrons+l-spins schemes, spins are treated classically whereas electrons quantum mechanically [30, 31, 32, 33, 34, 35, 36, 37, 38], which makes computations viable for fairly large samples. Such mixed quantum-classical strategy is reliable for skyrmion sizes of hundreds of lattice constants and/or large spin ($S\gtrsim 2$), where the role of quantum fluctuations are expected to be small. However, as experiments motivate, for nanometer-scale skyrmions [23] (and for $S\lesssim 1$) a quantum description [41, 42, 43, 39, 44] is in order, since outcomes from classical-spins can be at variance with quantum results [45]. Moreover, nanometer-scale skyrmions, due to their reduced spatial extent, exhibit pronounced susceptibility to boundary conditions, finite-size constraints, external field perturbations, and spin-transfer-induced phenomena such as spin slippage, all of which can critically influence the robustness of their topological protection.
@@
For spin-only quantum descriptions, calculations are usually via exact diagonalization (ED) and Tensor Networks (TN) [46, 47, 48, 49] (for a recent analytical study, see [44]). Another way to proceed is to treat the quantum skyrmion Hamiltonian at the mean-field level. [50]. With this simplified yet computationally efficient approach, both equilibrium and real-time dynamics become tractable even for very large systems. The advantage over a fully classical approach lies in the inclusion of local spin quantum fluctuations, though the accuracy of the mean-field description compared to fully quantum skyrmion solutions remains essentially unexamined in general.
@@
When both localized spins and itinerant electrons are explicitly considered for skyrmion-type Hamiltonians within a full quantum description, exact results are so far limited to very small systems, and primarily for benchmark uses; furthermore, treatments where the $l$-spin subsystem is treated at the mean-field level are also lacking. It is thus highly desirable (in fact necessary in several cases) to have a framework including i-electrons and l-spin textures, and in the presence of reservoirs/leads, that at the same time does not neglect the quantum effects on nanoskyrmions.
@@
In a lattice-site representation, a common way to consider such interaction is to have at each lattice site $R$ a local Kondo-like exchange coupling between i-electrons and l-spins , i.e. $g\sum_{R}\hat{\bf{s}}_{R}^{i}\cdot\hat{\bf{S}}_{R}^{l}$. However, to proceed, approximations are needed, since in essence one is then dealing with a Kondo-lattice-like type of problem, further complicated by Heisenberg and DMI interactions for the l-spins  [^1] In [39], a two-band, real space DMFT approach to skyrmions was considered, with the dynamics treated in the linear response regime, and in the absence of electronic reservoirs..
@@
In this work, we propose a practically viable quantum treatment for both skyrmions and electrons, in- and out-of-equilibrium. The approach uses an important level of approximation, namely a mean-field treatment of the Kondo-like exchange. Nevertheless, it permits to characterize intrinsically quantum physical quantities and gives access to the full nonequilibrium dynamics (i.e., well beyond the linear-response regime), with electronic reservoirs included in the formulation. From now on, we will refer to this scheme as QQmf, meaning that the we describe at the mean-field level the interaction between the two subsystems (i-electrons and l-spins) which are treated individually at the quantum level. Furthermore, we denote by QC the scheme with quantum i-electrons and classical l-spins, and by Wmf the scheme with quantum i-electrons and quantum mean-field l-spins.
@@
In the rest of the paper we exclusively consider the case of Neel-type Skyrmions, in situations where the Kondo-like coupling is ferromagnetic in character; the antiferromagnetic case is only briefly addressed in the supplemental material, (SM).
@@
The main outcomes of our study are: i) exact numerical benchmarks show that our method has higher accuracy over the QC approach; this statement is supported by results for both very small and larger systems, reported in the paper and in the SM. ii) in many of the benchmark situations examined, the Wmf descriptions improves over the QC one, and is very close to the QQmf result. Still, for some physical quantities, noticeable discrepancy remains between the Wmf and the QQmf, in both static and nonequilibrium regimes. iii) the itinerant electrons significantly influence a quantum skyrmion texture and its entanglement, quantum-chirality, and quantum-susceptibility properties; iv) in the dynamical regime following the application of a bias, quantum “nanomerons” and bond currents dynamically affect each other in a highly non trivial way at the atomistic lattice-site level. These features are not accessible to spin-only descriptions and/or are described differently within QC and Wmf schemes.
@@
System and Hamiltonian. As “proof of concept” of the method, we consider three setups of increasing complexity, shown in Fig. 1. By denoting as $C$ ($L_{E}$) the regions where the l-spins (i-electrons) reside, we have i) both l-spins and i-electrons in the same finite lattice $C\equiv L_{E}$; ii) l-spins in $C$, and i-electrons in an enlarged finite lattice $L_{E}=L+C+R$, where region $L$ ($R$) is coupled to $C$ from the left (right); iii) same as ii), but with semi-infinite $L$ and $R$ regions. We describe i-iii) via a Kondo lattice Hamiltonian plus Heisenberg exchange and DMI, i.e. $\hat{H}=\hat{H}^{s}+\hat{H}^{d}+\hat{H}^{sd}$ 222 This Hamiltonian can be derived from a two-band electron-only model, with electrons highly localized and correlated in one band (e.g., of the $d$ or $f$ type), and itinerant in the other band (e.g., of $s$ or $p$ type) [35], with
@@
Here, $\hat{H}^{s}$ is the i-electrons part, with hopping parameter $t_{h}$ between nearest-neighbour (n.n.) sites, and $\hat{c}_{i\sigma}^{\dagger}$ creating an electron with spin projection $\sigma$ at site $i$. The l-spins texture is described by $\hat{H}^{d}$, where $\mathbf{\hat{S}}_{i}$ is the spin operator at site $i$, and $J$, $\mathbf{B}$, and ${\bf D}$ respectively denote the Heisenberg exchange, the external magnetic field, and the Neel-type DMI, namely ${\bf D}_{ij}=D\frac{{\bf r}_{i}-{\bf r}_{j}}{|{\bf r}_{i}-{\bf r}_{j}|}\wedge\hat{\bf z}$. Finally, $\hat{H}^{sd}$ is the Kondo-like exchange term , with $\mathbf{\hat{s}}_{i}=(1/2)\sum_{\tau\tau^{\prime}}\hat{c}^{\dagger}_{i\tau}\boldsymbol{\sigma}_{\tau\tau^{\prime}}\hat{c}_{i\tau^{\prime}}$ and $\boldsymbol{\sigma}\equiv(\sigma^{x},\sigma^{y},\sigma^{z})$ the vector of Pauli matrices. In $\hat{H}^{s}$, there is no spin-orbit interaction (SOI). Usually small, SOI can sometimes affect skyrmion stabilisation [53, 40, 54, 36]), but it is not considered here to limit the number of model parameters.
Figure 1: Sketch of the setups considered, where i-electrons and l-spins interact via local Kondo-like exchange of strength $g$.
Figure 2: Relative average error $\eta$ (Eq. 4) between of the proposed QQmf scheme with respect to the full TN solution for a 2$\times$2 (left) and a 3$\times$3 (right) plaquette. Results for $\eta$ are also shown for quantum electron+classical spin (QC) solution and for the case where the localized spins are treated at the quantum mean-field level (Wmf). The parameters used are $J\!=\!0.1,B^{z}\!=\!0.1,t_{h}\!=\!1$. For the case $J=1$, see main text and the SM.
Approach and method of solution. We will treat $\hat{H}^{sd}$ at the mean-field level. Hence, by defining $\hat{V}^{s}=\sum_{i\in C}\langle\hat{\mathbf{S}}_{i}\rangle\cdot\hat{\mathbf{s}}_{i}$ and $\hat{V}^{d}=\sum_{i\in C}\langle\hat{\mathbf{s}}_{i}\rangle\cdot\hat{\mathbf{S}}_{i}$, we have $\hat{H}\rightarrow(\hat{H}^{s}_{MF},\hat{H}^{d}_{MF})$, with
$\hat{H}^{s/d}_{MF}=\hat{H}^{s/d}-g\hat{V}^{s/d}.$ (2)
@@
Isolated/enlarged system.- Starting with $\langle\mathbf{s}_{i}\rangle=0$, we use a Matrix Product State (MPS) algorithm [55, 56, 57] from the ITensor library [58] to find the ground state of $\hat{H}^{d}_{MF}$ and update the averages $\langle\mathbf{S}_{i}\rangle$. These then enter as parameters when solving via ED for $\hat{H}^{s}_{MF}$, while reproducing $\langle\mathbf{s}_{i}\rangle$. As the iterations converge, the mean-field ground state is reached (see the supplementary material, SM, for details). For the enlarged case, the procedure is carried out with $\epsilon_{i\sigma}=0$ at $t=0$). The time evolution for the enlarged system is discussed later in the paper.
Open system.- Starting with the ground state of the isolated region $C$, the tunnelling matrix elements between $C$ and the leads $R$, $L$ are switched on in time adiabatically up to a final value, so that $L+C+R$ reaches a steady state (chosen as the initial/ground state of the open system). In the process, the quantum l-spins evolve in time via time-evolution block-decimation (TEBD) [59], and the l-electrons via the nonequilibrium Green’s function (NEGF) approach [60, 61, 62, 63] in the equal-time formulation. In this scheme, the key quantity is the one-particle density matrix $\rho$ for region C, related to the full one-particle NEGF $G(t,t^{\prime})$ via $\rho(t)=-iG(t,t^{+})$. With leads, $\rho$ obeys
@@
Here, $h(t)$ is the (possibly time-dependent) single particle term and $I_{col.}$ is the collision integral [64], that in general describes interaction-induced correlation effects and lead-induced embedding effects. The i-electrons do not interact in our model, i.e. $I_{col.}$ has only the embedding part. We also i) use the so-called wide-band limit (WBL), where the hopping term $t_{l}$ in $L,R$ tends to infinity, whilst $\gamma=t_{l}^{2}/t_{h}$ remain fixed ($\gamma$ thus sets the magnitude of the $L$-$C$ and $R$-$C$ couplings); ii) further approximate the WBL by simplifying the calculation of $I_{col.}$. Steps i,ii) are a good trade-off between reduced computation time and accuracy [38] (see the SM).
@@
Mean-field and classical schemes for the localized spins.- As terms of comparison to the QQmf method, we considered i) quantum mean-field and ii) classical descriptions of the spin texture, while keeping the electrons and the Kondo-like coupling treated as in the QQmf scheme. Description i) is based on the Weiss-like mean-field (Wmf) approach from [50] and briefly summarized in the SM; details of ii) can be found in e.g. [30, 38].
@@
Figure 3: Ground state results for a) and b): an isolated system with $D=0.2,J=0.1,B^{z}=0.1,t_{h}=1,g=2$, and c): an open system with $D=1,J=0.2,B^{z}=0.5,t_{h}=1,g=1$. a) Spin expectation-value (left) and concurrences $\mathcal{C}$ (right) heat maps in a $21\times 21$ isolated rhombus cluster, with l-spins (i-electrons) chiralities $Q_{\text{sp}}=-6.37$ ($Q_{\text{el}}=-0.03$). In the spin map, the top (bottom) half shows the expectation value $\langle S^{x,y,z}\rangle$ ($\langle s^{x,y,z}\rangle$) of the l-spins (the i-electrons). The values of the $z$- and $xy$ spin-components are indicated by colours and arrows, respectively. In the concurrence map, the value of $\mathcal{C}$ for n.n. (upper half) and next nearest neighbour (2nn, lower half) sites is represented by the bond colour. The missing half of each map is recovered using rhombus symmetry. b) Logarithm of the static structure factors $\ln S^{zz}$ (left) and $\ln S^{+-}$ (right) of the l-spins. c) $\langle S^{x,y,z}\rangle$ (left half) and $\langle s^{x,y,z}\rangle$ for a $6\times 6$ square cluster contacted to leads, with parameters $D=1,J=0.2,B^{z}=0.5,t_{h}=1,g=1$.
@@
Figure 4: Nonequilibrium dynamics of a rhombus cluster $C$ of $4\times 10$ sites with an enlarged electron region $L+C+R$ of $4\times(100+10+100)$ sites, for parameters $D\!=\!0.2,J\!=\!0.05,B^{z}\!=\!0.06,t_{h}\!=\!1,g\!=\!2$. The dynamics is induced by a spin-up polarized bias $\epsilon_{L}=-\epsilon_{R}$ in the enlarged region, switched-on at $t\!=\!0$ and ramped to maximum strength 0.5 at $\!t=\!10$. a) Spin-up current $I_{\langle ij\rangle}$ along the bonds $\langle ij\rangle$ labeled by A-F (sketch on the right). b) Time snapshots of the spin expectation values $\langle S^{x,y,z}(t)\rangle$ in $C$, $z$- and $xy$-spin components indicated by colours and arrows, respectively. The current-driven motion of one meron is highlighted by the green semicircle.
@@
Exact vs mean-field Kondo-like exchange, and quantum vs classical spins. To see if a quantum treatment of l-spins+i-electrons (albeit within a mean-field account of the Kondo-like exchange) is beneficial, we consider 2$\times$2 and 3$\times$3 square plaquettes (respectively with 4 spins + 4 electrons and 9 spins + 9 electrons) described by Eq. (1), and with open boundary conditions (OBC). Via the TN method, we determine full quantum ground state of these systems, which is then used as reference to assess the scope of the QQmf and QC descriptions. We do this via the quantity (A = QQmf or QC)
$\eta_{\text{A}}=\frac{1}{N}\sum_{i}\frac{|\langle{\hat{\bm{S}}_{i}}\rangle_{\text{A}}-\langle{\hat{\bm{S}}_{i}}\rangle_{TN}|}{|\langle{\hat{\bm{S}}_{i}}\rangle_{TN}|},$ (4)
@@
The indication from panels a),b) of Fig. 2 is that QQmf and QC give similar and overall satisfactory results for the 2$\times$2 system, although the QC description is slightly inferior in a subregion of the heat-map. On the other hand, for the 3$\times$3 plaquette (panels c,d), the QQMF solution is markedly better, especially in the region corresponding to large $g$ and $D$ (for reference, the QQmf treatment is exact for $g=0$, as reflected in the heat maps). While providing limited evidence, the results of Fig. 2 could also hint that the difference in quality of the QQmf and QC treatments may increase for larger system size, thus lending support to the choice of parameter values considered in the rest of the paper.
@@
Besides the results of Fig. 2, in the SM we report several other comparisons in terms of the 2$\times$2 and 3$\times$3 systems, and for both FM ($g>0$) and AF ($g<0$) couplings between i-electrons and l-spins. The overall picture emerging from Fig. 2 and the additional testing is that for the FM case there is an appreciable region of the $g,D$ values where the QQmf approach performs better than the QC treatment, whilst at small $g,D$ values QQmf and QC retain a similar (good) performance. This behavior occurs for either small or large values of the FM coupling among $\it l$-spins. On the other hand, in the AF case, both QQmf and QC do not perform well already for rather moderate $g,D$ values, due to the inadequacy of the mean-field treatment in the Kondo regime.

### v4

@@ Abstract
A quantum description is given of nanoskyrmions in two-dimensional textures with localized spins and itinerant electrons, isolated or coupled to leads, in or out of equilibrium. The spin-electron exchange is treated at the mean-field level, while tensor networks and exact diagonalization or nonequilibrium Green’s functions are used for localized spins and itinerant electrons. We motivate our scheme via exact and mean-field benchmarks, then show by several examples that itinerant electrons distinctly affect the properties of quantum nanoskyrmions. Finally, we mention lines of future work and improvement of the approach.
@@
Introduction. Magnetic skyrmions are topologically robust, long-lived spin textures which can be manipulated by ultralow currents [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11], and of great potential for applications in spintronics and quantum computing [12, 13, 14, 15, 16]. Skyrmions originate from the interplay of different types of magnetic interactions in solids [17], such as Heisenberg exchange [18], Dzyaloshinskii-Moriya interaction (DMI) [19, 20], and anisotropy [21, 22], to mention a few.
@@
Theoretical descriptions of skyrmions are often based on classical spin-only models [23, 24, 25, 26], treated via the Landau-Lifshitz-Gilbert [27, 28] and Thiele equations [29]. These formulations, where itinerant electrons (i electrons) enter implicitly or as a source of renormalization of the classical spin dynamics, are in many cases adequate; however, describing explicitly the interaction (via Kondo-like exchange) between i electrons and localized spins (l spins) can also be important [30, 31, 32, 33, 34, 35, 36, 37, 38, 39], to, e.g., address interface phenomena like current driving and optical generation of nanoskyrmions [40].
@@
In most of currently available i electrons+l spins schemes, spins are treated classically whereas electrons quantum mechanically [30, 31, 32, 33, 34, 35, 36, 37, 38], which makes computations viable for fairly large samples. Such mixed quantum-classical strategy is reliable for skyrmion sizes of hundreds of lattice constants and/or large spin ($S\gtrsim 2$), where the role of quantum fluctuations is expected to be small. However, in small systems, theoretical treatments based on classical and quantum spins can give different results (see e.g. Refs. [41, 42]), suggesting that, to characterize experiments on nanometer-scale [23] skyrmions, a quantum description of spins [43, 44, 42, 39, 45] is better suited. Moreover, nanometer-scale skyrmions, due to their reduced spatial extent, exhibit pronounced susceptibility to boundary conditions, finite-size constraints, external field perturbations, and spin-transfer-induced phenomena such as spin slippage, all of which can critically influence the robustness of their topological protection.
@@
For spin-only quantum descriptions, calculations are usually via exact diagonalization (ED) and tensor networks (TNs) [46, 47, 48, 49] (for a recent analytical study, see [45]). Another way to proceed is to treat the quantum skyrmion Hamiltonian at the mean-field level [50]. With this simplified yet computationally efficient approach, both equilibrium and real-time dynamics become tractable even for very large systems. The advantage over a fully classical approach lies in the inclusion of local spin quantum fluctuations, though the accuracy of the mean-field description compared to fully quantum skyrmion solutions remains essentially unexamined in general.
@@
When both localized spins and itinerant electrons are explicitly considered for skyrmion-type Hamiltonians within a full quantum description, exact results are so far limited to very small systems, and primarily for benchmark uses; furthermore, treatments where the $l$ spin subsystem is treated at the mean-field level are also lacking. It is thus highly desirable (in fact, necessary in several cases) to have a framework including i electrons and l spin textures, and in the presence of reservoirs or leads, that at the same time does not neglect the quantum effects on nanoskyrmions.
@@
In a lattice-site representation, a common way to consider such interaction is to have at each lattice site $R$ a local Kondo-like exchange coupling between i electrons and l spins , i.e. $g\sum_{R}\hat{\bf{s}}_{R}^{i}\cdot\hat{\bf{S}}_{R}^{l}$. However, to proceed, approximations are needed, since in essence one is then dealing with a Kondo-lattice-like type of problem, further complicated by Heisenberg and DMI interactions for the l spins  [^1] In [39], a two-band, real space DMFT approach to skyrmions was considered, with the dynamics treated in the linear response regime, and in the absence of electronic reservoirs..
@@
In this work, we propose a practically viable quantum treatment for both skyrmions and electrons, in and out of equilibrium. The approach uses an important level of approximation, namely a mean-field treatment of the Kondo-like exchange. Nevertheless, it permits one to characterize intrinsically quantum physical quantities and gives access to the full nonequilibrium dynamics (i.e., well beyond the linear-response regime), with electronic reservoirs included in the formulation. From now on, we will refer to this scheme as QQmf, meaning that the we describe at the mean-field level the interaction between the two subsystems (i electrons and l spins) which are treated individually at the quantum level. Furthermore, we denote by QC the scheme with quantum i electrons and classical l spins, and by Wmf the scheme with quantum i electrons and quantum mean-field l spins.
@@
In the rest of the Letter we exclusively consider the case of Néel-type skyrmions, in situations where the Kondo-like coupling is ferromagnetic in character; the antiferromagnetic case is only briefly addressed in the Supplemental Material (SM)  [^2] See Supplemental Material at [URL to be inserted] for a brief discussion of antiferromagnetic Kondo coupling. The Supplemental Material also includes computational details and additional benchmark comparisons..
@@
The main outcomes of our study are as follows: (i) Exact numerical benchmarks show that our method has higher accuracy over the QC approach; this statement is supported by results for both very small and larger systems, reported in the Letter and in the SM. (ii) In many of the benchmark situations examined, the Wmf description improves over the QC one, and is very close to the QQmf result. Still, for some physical quantities, a noticeable discrepancy remains between the Wmf and the QQmf, in both static and nonequilibrium regimes. (iii) The itinerant electrons significantly influence a quantum skyrmion texture and its entanglement, quantum chirality, and quantum susceptibility properties. (iv) In the dynamical regime following the application of a bias, quantum “nanomerons” and bond currents dynamically affect each other in a highly nontrivial way at the atomistic lattice-site level. These features are not accessible to spin-only descriptions and/or are described differently within QC and Wmf schemes.
@@
System and Hamiltonian. As “proof of concept” of the method, we consider three setups of increasing complexity, shown in Fig. 1. By denoting as $C$ ($L_{E}$) the regions where the l spins (i electrons) reside, we have (i) both l spins and i electrons in the same finite lattice $C\equiv L_{E}$; (ii) l spins in $C$, and i electrons in an enlarged finite lattice $L_{E}=L+C+R$, where region $L$ ($R$) is coupled to $C$ from the left (right); (iii) same as (ii), but with semi-infinite $L$ and $R$ regions. We describe (i–iii) via a Kondo lattice Hamiltonian plus Heisenberg exchange and DMI, i.e., $\hat{H}=\hat{H}^{s}+\hat{H}^{d}+\hat{H}^{sd}$ 333 This Hamiltonian can be derived from a two-band electron-only model, with electrons highly localized and correlated in one band (e.g., of the $d$ or $f$ type), and itinerant in the other band (e.g., of $s$ or $p$ type) [35], with
@@
Here, $\hat{H}^{s}$ is the i electrons part, with hopping parameter $t_{h}$ between nearest-neighbor (n.n.) sites, and $\hat{c}_{i\sigma}^{\dagger}$ creating an electron with spin projection $\sigma$ at site $i$. The l spins texture is described by $\hat{H}^{d}$, where $\mathbf{\hat{S}}_{i}$ is the spin operator at site $i$, and $J$, $\mathbf{B}$, and ${\bf D}$, respectively, denote the Heisenberg exchange, the external magnetic field, and the Néel-type DMI, namely, ${\bf D}_{ij}=D\frac{{\bf r}_{i}-{\bf r}_{j}}{|{\bf r}_{i}-{\bf r}_{j}|}\wedge\hat{\bf z}$. Finally, $\hat{H}^{sd}$ is the Kondo-like exchange term , with $\mathbf{\hat{s}}_{i}=(1/2)\sum_{\tau\tau^{\prime}}\hat{c}^{\dagger}_{i\tau}\boldsymbol{\sigma}_{\tau\tau^{\prime}}\hat{c}_{i\tau^{\prime}}$ and $\boldsymbol{\sigma}\equiv(\sigma^{x},\sigma^{y},\sigma^{z})$ the vector of Pauli matrices. In $\hat{H}^{s}$, there is no spin-orbit interaction (SOI). Usually small, SOI can sometimes affect skyrmion stabilization [54, 40, 55, 36]), but it is not considered here to limit the number of model parameters.
Figure 1: Sketch of the setups considered, where i electrons and l spins interact via local Kondo-like exchange of strength $g$.
Figure 2: Relative average error $\eta$ [Eq. 4] between the proposed QQmf scheme with respect to the full TN solution for a 2$\times$2 (left) and a 3$\times$3 (right) plaquette. Results for $\eta$ are also shown for quantum electron+classical spin (QC) solution and for the case where the localized spins are treated at the quantum mean-field level (Wmf). The parameters used are $J\!=\!0.1,B^{z}\!=\!0.1,t_{h}\!=\!1$. For the case $J=1$, see the main text and the SM.
Approach and method of solution. We will treat $\hat{H}^{sd}$ at the mean-field level. Hence, by defining $\hat{V}^{s}=\sum_{i\in C}\langle\hat{\mathbf{S}}_{i}\rangle\cdot\hat{\mathbf{s}}_{i}$ and $\hat{V}^{d}=\sum_{i\in C}\langle\hat{\mathbf{s}}_{i}\rangle\cdot\hat{\mathbf{S}}_{i}$, we have $\hat{H}\rightarrow(\hat{H}^{s}_{\text{MF}},\hat{H}^{d}_{\text{MF}})$, with
$\hat{H}^{s/d}_{\text{MF}}=\hat{H}^{s/d}-g\hat{V}^{s/d}.$ (2)
@@
Isolated or enlarged system. Starting with $\langle\mathbf{s}_{i}\rangle=0$, we use a matrix product state (MPS) algorithm [56, 57, 58] from the ITensor library [59] to find the ground state of $\hat{H}^{d}_{\text{MF}}$ and update the averages $\langle\mathbf{S}_{i}\rangle$. These then enter as parameters when solving via ED for $\hat{H}^{s}_{\text{MF}}$, while reproducing $\langle\mathbf{s}_{i}\rangle$. As the iterations converge, the mean-field ground state is reached (see the Supplemental Material (SM) for details). For the enlarged case, the procedure is carried out with $\epsilon_{i\sigma}=0$ at $t=0$. The time evolution for the enlarged system is discussed later in the Letter.
Open system. Starting with the ground state of the isolated region $C$, the tunneling matrix elements between $C$ and the leads $R$, $L$ are switched on in time adiabatically up to a final value, so that $L+C+R$ reaches a steady state (chosen as the initial or ground state of the open system). In the process, the quantum l spins evolve in time via time-evolution block-decimation (TEBD) [60], and the l-electrons via the nonequilibrium Green’s function (NEGF) approach [61, 62, 63, 64] in the equal-time formulation. In this scheme, the key quantity is the one-particle density matrix $\rho$ for region C, related to the full one-particle NEGF $G(t,t^{\prime})$ via $\rho(t)=-iG(t,t^{+})$. With leads, $\rho$ obeys
@@
Here, $h(t)$ is the (possibly time-dependent) single-particle term and $I_{\text{col.}}$ is the collision integral [65], that in general describes interaction-induced correlation effects and lead-induced embedding effects. The i electrons do not interact in our model; i.e., $I_{\text{col.}}$ has only the embedding part. We also (i) use the so-called wide-band limit (WBL), where the hopping term $t_{l}$ in $L,R$ tends to infinity, while $\gamma=t_{l}^{2}/t_{h}$ remains fixed ($\gamma$ thus sets the magnitude of the $L$-$C$ and $R$-$C$ couplings), and (ii) further approximate the WBL by simplifying the calculation of $I_{\text{col.}}$. Steps (i,ii) are a good trade-off between reduced computation time and accuracy [38] (see the SM).
@@
Mean-field and classical schemes for the localized spins. As terms of comparison to the QQmf method, we considered (i) quantum mean-field and (ii) classical descriptions of the spin texture, while keeping the electrons and the Kondo-like coupling treated as in the QQmf scheme. Description (i) is based on the Weiss-like mean-field (Wmf) approach taken from [50] and briefly summarized in the SM; details of (ii) can be found in, e.g., [30, 38].
@@
Figure 3: Ground state results for (a), (b) an isolated system with $D=0.2,J=0.1,B^{z}=0.1,t_{h}=1,g=2$, and (c): an open system with $D=1,J=0.2,B^{z}=0.5,t_{h}=1,g=1$. (a) Spin expectation value (left) and concurrences $\mathcal{C}$ (right) heat maps in a $21\times 21$ isolated rhombus cluster, with l spin (i electron) chiralities $Q_{\text{sp}}=-6.37$ ($Q_{\text{el}}=-0.03$). In the spin map, the top (bottom) half shows the expectation value $\langle S^{x,y,z}\rangle$ ($\langle s^{x,y,z}\rangle$) of the l spins (the ielectrons). The values of the $z$- and $xy$ spin-components are indicated by colours and arrows, respectively. In the concurrence map, the value of $\mathcal{C}$ for n.n. (upper half) and next-nearest-neighbour (2n.n., lower half) sites is represented by the bond color. The missing half of each map is recovered using rhombus symmetry. (b) Logarithm of the static structure factors $\ln S^{zz}$ (left) and $\ln S^{+-}$ (right) of the l spins. (c) $\langle S^{x,y,z}\rangle$ (left half) and $\langle s^{x,y,z}\rangle$ for a $6\times 6$ square cluster contacted to leads, with parameters $D=1,J=0.2,B^{z}=0.5,t_{h}=1,g=1$.
@@
Figure 4: Nonequilibrium dynamics of a rhombus cluster $C$ of $4\times 10$ sites with an enlarged electron region $L+C+R$ of $4\times(100+10+100)$ sites, for parameters $D\!=\!0.2,J\!=\!0.05,B^{z}\!=\!0.06,t_{h}\!=\!1,g\!=\!2$. The dynamics is induced by a spin-up polarized bias $\epsilon_{L}=-\epsilon_{R}$ in the enlarged region, switched on at $t\!=\!0$ and ramped to maximum strength 0.5 at $\!t=\!10$. (a) Spin-up current $I_{\langle ij\rangle}$ along the bonds $\langle ij\rangle$ labeled by A–F (sketch on the right). (b) Time snapshots of the spin expectation values $\langle S^{x,y,z}(t)\rangle$ in $C$, $z$- and $xy$-spin components indicated by colors and arrows, respectively. The current-driven motion of one meron is highlighted by the green semicircle.
@@
Exact versus mean-field Kondo-like exchange, and quantum versus classical spins. To see if a quantum treatment of l spins+i electrons (albeit within a mean-field account of the Kondo-like exchange) is beneficial, we consider 2$\times$2 and 3$\times$3 square plaquettes (respectively, with 4 spins + 4 electrons and 9 spins + 9 electrons) described by Eq. (1), and with open boundary conditions (OBC). Via the TN method, we determine the full quantum ground state of these systems, which is then used as a reference to assess the scope of the QQmf and QC descriptions. We do this via the quantity (A = QQmf or QC)
$\eta_{\text{A}}=\frac{1}{N}\sum_{i}\frac{|\langle{\hat{\bm{S}}_{i}}\rangle_{\text{A}}-\langle{\hat{\bm{S}}_{i}}\rangle_{\text{TN}}|}{|\langle{\hat{\bm{S}}_{i}}\rangle_{\text{TN}}|},$ (4)
@@
The indication from panels (a,b) of Fig. 2 is that QQmf and QC give similar and overall satisfactory results for the 2$\times$2 system, although the QC description is slightly inferior in a subregion of the heat-map. On the other hand, for the 3$\times$3 plaquette [panels (c,d)], the QQMF solution is markedly better, especially in the region corresponding to large $g$ and $D$ (for reference, the QQmf treatment is exact for $g=0$, as reflected in the heat maps). While providing limited evidence, the results of Fig. 2 could also hint that the difference in quality of the QQmf and QC treatments may increase for larger system size, thus lending support to the choice of parameter values considered in the rest of the Letter.
@@
Besides the results of Fig. 2, in the SM we report several other comparisons in terms of the 2$\times$2 and 3$\times$3 systems, and for both ferromagnetic (FM) ($g>0$) and antiferromagnetic (AF) ($g<0$) couplings between i electrons and l spins. The overall picture emerging from Fig. 2 and the additional testing is that for the FM case there is an appreciable region of the $g,D$ values where the QQmf approach performs better than the QC treatment, while at small $g,D$ values QQmf and QC retain a similar (good) performance. This behavior occurs for either small or large values of the FM coupling among $\it l$ spins. On the other hand, in the AF case, both QQmf and QC do not perform well already for rather moderate $g,D$ values, due to the inadequacy of the mean-field treatment in the Kondo regime.